### v3

# PropNet: a White-Box and Human-Like Network for Sentence Representation

Abstract
Transformer-based embedding methods have dominated the field of sentence representation in recent years. Although they have achieved remarkable performance on NLP missions, such as semantic textual similarity (STS) tasks, their black-box nature and large-data-driven training style have raised concerns, including issues related to bias, trust, and safety. Many efforts have been made to improve the interpretability of embedding models, but these problems have not been fundamentally resolved. To achieve inherent interpretability, we propose a purely white-box and human-like sentence representation network, PropNet. Inspired by findings from cognitive science, PropNet constructs a hierarchical network based on the propositions contained in a sentence. While experiments indicate that PropNet has a significant gap compared to state-of-the-art (SOTA) embedding models in STS tasks, case studies reveal substantial room for improvement. Additionally, PropNet enables us to analyze and understand the human cognitive processes underlying STS benchmarks.
1 Introduction
Sentence representation research has been carried out for over half a century, evolving from regular expressions [23], statistical encoding methods [27; 36; 50; 32] to neural network embeddings [42; 10; 35; 34; 20; 47; 22; 39]. In recent years, embedding models have dominated the research community, and also have achieved significant success in industry applications. However, their black-box nature and reliance on large-scale data-driven training raise concerns about bias [11; 14; 13; 44], trust [48; 49], safety [8; 46] and other issues [9]. Researchers have attempted to improve the interpretability of embedding models [30; 40] and remove their biases [26; 25], but these problems remain fundamentally unsolved. Moreover, in the process of training models, massive amounts of manually labeled data have been generated. However, researchers typically focus solely on the training value of these labeled data, neglecting the cognitive values that exist within them. These cognitive values hold significant potential for enhancing human self-understanding.
In the cognitive science community, the question of how our brains represent a sentence has drawn a great deal of research interest [37; 7; 6; 21; 24; 29; 31]. When humans read and comprehend a sentence, they decompose it into propositions. A proposition is a simple sentence with no more than one verb. The meanings of these propositions are stored in long-term memory as a hierarchical network, which can be recalled for use in downstream tasks [38; 18]. The brain’s verbal short-term memory possesses a specialized subsystem. This subsystem is responsible for splitting propositions and handling the syntactic features of each proposition as the eyes sweep across the text [15]. Moreover, the cognitive load in sentence comprehension is primarily attributed to syntactic complexity, such as the number of verbs, rather than superficial length, such as the addition of adjectives and other modifiers [24]. This implies proposition is a processing unit for sentence comprehension. Inspired by these findings, in this work, we propose a purely white-box and human-like sentence representation network named PropNet.
Method No Feature Engineering Semantic Preservation WhiteBox HumanLike
Regular Expressions $\times$ $\times$ $\checkmark$ $\times$
Statistical Encoding $\checkmark$ $\times$ $\checkmark$ $\times$
Neural Network Embedding $\checkmark$ $\checkmark$ $\times$ $\times$
CISM $\checkmark$ $\times$ $\checkmark$ $\checkmark$
PropNet $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison of Sentence Representation Methods
PropNet constructs a hierarchical network based on the propositions comprising a sentence. The generation of PropNet involves four phases: splitting, parsing, representing and merging. The first two phases employ a mechanism analogous to short-term memory, splitting a sentence into propositions and parsing key tokens from each proposition according to specific dimensions. The dimensions include Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, and Source. The last two phases function akin to long-term memory, representing a proposition as a hierarchical network and merging all networks into a unified network to express the complete meaning of the original sentence. Figure 3 provides an example of PropNet.
PropNet is purely white-box since it is completely interpretable and transparent. It is also human-like because it has a hierarchy network structure to save the meaning in a sentence. With respect to semantics, unlike embedding models which focus more on the relationships between words and their contextual information, PropNet focuses on the true happenings in the world as expressed by a sentence. It achieves this by analyzing the core components of propositions, particularly the eight dimensions related to verbs. The principles behind PropNet’s design are very natural, without specially crafted artificial features. The comparison of PropNet and other representation methods is summarized in Table 1.
In semantic textual similarity (STS) tasks, given a pair of sentences, humans are required to assign a similarity score according to specific instructions. Cognitive processes play a key role in bridging sentences/instructions and annotation results. PropNet enables us to study this cognitive process. Experiments show that differences in actions, subjects or objects are a sufficient condition for humans to confirm an inequivalent proposition pair.
We compare PropNet with representative embedding methods on STS tasks. Although experiments imply that PropNet has a large gap compared to state-of-the-art (SOTA) embedding models, case studies reveal significant potential for enhancement.
2 Relative Work
Text representation approaches could date back to rule-based regular expressions for search engines and question-answering systems [23]. Feature engineering is implemented heavily to develop expressions for pattern matching. This approach is inherently explainable, as the expressions consist of readable strings. However, it is pretty rigid and shallow, without a deep understanding of semantics.
Statistical encoding methods compute word frequencies and represent text as discrete or continuous vectors. Typical methods include One Hot Embedding (OHE), Bag of Words (BOW) [27], N-grams [36], Normalized Term Frequency (NTF) [50] and TF-IDF [32]. These methods are straightforward to implement since no feature engineering is required. They are also explainable, as each number in the representation vectors has a clear meaning. However, since word frequency cannot handle semantic issues, like synonyms and antonyms, or word relationships, these methods only provide shallow semantic analysis.
Neural network embedding approaches represent words as vectors in a continuous vector space. Representative methods include Word2Vec [42], FastText [10; 35; 34], Glove [45], BERT [20], SBERT [47], SimCSE [22] and AoE [39]. They train neural networks on an extremely large corpus, using different Language Modeling (LM) tasks as the objective. Since embeddings are automatically computed by backpropagation, feature engineering is completely eliminated. In addition, the embeddings provide deep semantic understanding of the text, as they capture the semantic aspects of language due to LM tasks as the training goal.
Embedding methods are considered as black-box algorithms because the embedding vectors are generally unexplainable. This characteristic surreptitiously encodes social biases contained in the training corpus [11; 14; 13; 44], and also raises concerns about trust [48; 49], safety [8; 46] and other issues [9]. Numerous efforts have been made to improve interpretability [30; 40] and remove biases [26; 25], but these problems have not been fundamentally solved.
All the aforementioned methods are not human-like, as their representation formats are either strings or vectors. Cognitive-Inspired Syntactic Models (CISM), such as Tree Adjoining Grammar [33], Combinatory Categorial Grammar [51], and Dependency Grammar [28], are based on principles of human language comprehension and production. These models use interpretable tree structures to represent sentences but primarily focus on syntactic organization. PropNet, which employs Dependency Grammar for parsing, enhances semantic representation through instance nodes and stamp nodes. Unlike CISM, PropNet uses a network structure rather than a tree structure to represent sentences. Table 1 summarizes the characteristics of these methods.
3 Methods
Figure 1: The framework of building PropNet. The type of the input sentence is calculated before splitting. After merging, all networks are integrated into one network as the PropNet of the input sentence. Note that for types Prop0 and Prop1 the splitting and merging phases are omitted.
Clause Type Primary Rule Example
Conjunct token.dep_ == conj I am singing and playing a guitar. -> (I am singing, I be playing a guitar)
Clausal Complement token.dep_ == ccomp She thinks this is a good idea. -> (she thinks identifier_ccomp, this is a good idea)
Adnominal Clause token.dep_ == acl A cat sitting on sand looks up at the camera. -> (a cat looks up at the camera, cat be sitting on sand)
Prepositional Complement token.dep_ == pcomp He goes to school by taking a bus. -> (he goes to school by identifier_pcomp, he be taking a bus)
Adverbial Clause token.dep_ == advcl Although it is raining, we will go to a garden. -> (we will go to a garden, it is raining)
Open Clausal Complement token.dep_ == xcomp He is going to swim. -> (he is going to xcomp_identifier, he swim)
Subject Clause token.dep_ == csubj Adding aspirin to the water could kill the plant. -> (identifier_csubj could kill the plant, adding aspirin to the water)
Relative Clause token.dep_ == relcl They like the person who lives in the street. -> (They like the person, person lives in the street)
Table 2: Splitting rules for a Prop2 sentence according to its clause types. The condition token.pos_ in (VERB,AUX) is also required for all rules. We omit it to improve the clarity of the table. Note that token.pos_ and token.dep_ are the POS tagging and dependency parsing results from spaCy. The examples follow the format: Input Sentence -> (Main Proposition, Subordinate Proposition). The bold text in an example is the extracted verb token for the subordinate clause. Note that identifier_ccomp, identifier_pcomp, xcomp_identifier and identifier_csubj serve as placeholders for the relevant clauses.
To build a PropNet of a sentence, we first split the sentence into propositions (Section 3.1). Each proposition corresponds to a verb in the sentence. Then propositions are parsed to extract core tokens with respect to dimensions Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, and Source (Section 3.2). After that, each proposition is represented by a hierarchical network (Section 3.3). At last, all networks are merged into one network as the PropNet of the input sentence (Section 3.4). Figure 1 illustrates the framework of constructing PropNet. We introduce a method for comparing two PropNets, thereby unlocking the potential of the resulting PropNets for STS tasks (Section 3.5).
The following symbols are used to represent different types of sentences based on the number of verbs they contain:
  • Prop0: Sentences with 0 verbs, e.g., “A ball on the ground.”
  • Prop1: Sentences with 1 verb, e.g., “A boy is playing football.”
  • Prop2: Sentences with 2 verbs, e.g., “A boy is playing football and a girl is dancing.”
  • Prop3+: Sentences with more than 2 verbs, e.g., “A man is sitting in a chair, wearing a cloak, and holding a stick.”
Here the term "verb" encompasses linking verbs. For instance, in the sentences "a boy is in a park" or "a girl is happy," the word "is" is treated as a verb, and both sentences are categorized as Prop1. In the rest of this work, a proposition refers to a Prop1 or Prop0 sentence.
We introduce the following notations to classify sentence pairs based on sentence types:
  • P1- denotes the set of pairs where both sentences are Prop0 or Prop1.
  • P2 represents the set of pairs that contain at least one Prop2, excluding pairs with any Prop3. For instance, a Prop1 and a Prop2 sentence, or two Prop2 sentences.
  • P3+ indicates the set of pairs that includes at least one Prop3 sentence.
3.1 Splitting
Figure 2: Splitting a Prop3+ sentence and merging its proposition representations by backtracking. At each splitting step, the method for splitting Prop2 is called. The extracted proposition invokes the representation component to build a hierarchical network. During the upward traversal, the same strategies for merging P2—denoted by green, grey, and purple arrows—are employed to integrate all networks into a unified network. Note that identifier_advcl is a placeholder for an adverbial clause.
For Prop0 and Prop1, splitting is unnecessary. At this stage, the input sentence itself serves as the output proposition. For Prop2, we split the input sentence into two Prop1 propositions, a main proposition (main clause) and a subordinate proposition (subordinate clause). Table 2 presents the splitting rules for different clause types. These rules extract the verb token of the subordinate clause. Then the subordinate clause is generated by iterating over the verb token’s subtree provided by spaCy parsing tool  [^1] For more details, see the spaCy documentation: https://spacy.io/usage/linguistic-features/#pos-tagging. For instance, in the Prop2 sentence “She thinks this is a good idea”, the splitting rules identify “is” as the verb token for the subordinate clause. The token “is” owns a subtree “this is a good idea” by dependency parsing.
The generated subordinate clause is then replaced by a clause identifier in the input sentence for five clause types Clausal Complement, Prepositional Complement, Open Clausal Complement, Subject Clause and Adverbial Clause. Their identifiers are notated as identifier_ccomp, identifier_pcomp, xcomp_identifier, identifier_csubj and identifier_advcl, respectively. For other clause types, the subordinate clause is replaced by an empty string. When the replacement is completed, the rest of the input sentence becomes the main proposition. Table 2 presents examples for each clause type. Note that identifier_advcl is applied only when the clause is preceded by the word “to” that indicates the purpose of an action. Figure 2 provides an example for this specific situation.
Clause identifiers are crucial for merging proposition representations into a unified network. In the above example, the main proposition is “She thinks identifier_ccomp” where identifier_ccomp would point to “this is a good idea” in the merging phase.
If the subordinate clause is non-finite, the subject of the main clause is spliced to it, forming a complete subordinate proposition. For instance, in “she wants to dance”, the subordinate proposition “she dance” uses the subject of the main proposition “she wants xcomp_identifier”. If the subordinate clause is finite, no subject insertion is required since the clause already contains its own subject. The splitting rules for relative clause type are more intricate and are detailed in Appendix A.2.
Situation becomes significantly more complex for Prop3+. We employ a backtracking algorithm to decompose the input sentence into Prop1 propositions by constructing a binary tree, as shown in Figure 2. The splitting method for Prop2 is invoked during each splitting step to extract a subordinate proposition Prop1, serving as the leaf node of the binary tree. Then the leaf node invokes the representation component to construct a hierarchical network, which is explained in Section 3.3. The leaf nodes indicate the order of the propositions contained in the long input sentence from left to right. In Figure 2, the input sentence has an ordered proposition list [“A black and white dog is jumping in the air to identifier_advcl”, “dog catch a Frisbee”, “a man is getting identifier_ccomp”, “his boat clean”, “he wants a water voyage”] according to the leaf nodes from left to right. This order is necessary when we compare two Prop3+ sentences for similarity, which is explained in Section 3.5.
3.2 Parsing
Dimension Explanation Primary Rule Example
Action An action described by the verb in a proposition. token.pos_ is VERB The man is riding a horse.
Subject * An entity or concept that performs the action. token.dep_ is nsubj The man is riding a horse.
Object * An entity or concept that receives the action. token.dep_ in (dobj, nsubjpass) The man is riding a horse.
Where * The place where the action occurs, or the time when it occurs. preposition of token in “on”, “in”, “inside”, “outside”, “at”, …) The polar bear is sliding on the snow.
Aux_Obj * The instrument or tool that is used to perform the action. Or indirect object which is the recipient, beneficiary, or target of an action. preposition of token in (“with”, “by”, “about”, “as”, …) A person is cutting mushrooms with a knife.
Goal * The intended outcome or objective that the subject aims to achieve by performing the action. preposition of token in (“into”, “to”, “onto”, “towards”,…) A man pours oil into a pot.
Reason * The cause or motivation behind the action. preposition of token in (“for”, “due”) The teacher praised her for her excellent work.
Source * The origin or starting point from which the action arises. preposition of token in (“from”) She learned the news from her friend.
Attribute The appearance, state, nature, or features of an entity or a concept. token.dep_ in (amod, nmod, compound) A young girl is using sign language.
Part_of One entity is a component or piece of another larger entity. An entity token of another entity token; token.dep_ is poss Two men are packing suitcases into the trunk of a car.
Table 3: Dimensions of parsing a Prop1/Prop0 sentence. Dimensions Attribute and Part_of are designed for an entity or concept. Others are related to the action in a proposition. Primary rules of extracting tokens are explained with examples. Note that token.pos_ and token.dep_ are the POS tagging and dependency parsing results from spaCy. The bold text in an example is the extracted token for its dimension. Dimensions marked with a “*” also follow this parsing rule token.pos_ in (NOUN, PRON, PROPN, NUM) which is omited in the table for brevity.
This phase extracts key tokens from a proposition with respect to specific dimensions. The intuition of designing these dimensions is that we require they can describe the meaning of a single action that happens in the real world, which is the key to understand the meaning of a complex sentence consisting of a series of actions. Therefore, eight critical dimensions related to the verb of a proposition are considered: Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, Source. The extracted words under each dimension are either describing a physical entity like “car” or an abstractive concept like “brand”. For these words, two additional dimensions Attribute and Part_of are considered for the descriptions of them. Explanations for these dimensions and extracting primary rules  [^2] Because of the variety of language, even a Prop1 sentence can have an extremely flexible structure. Therefore, minor rules are handcrafted for extraction which can be seen in the codes. with examples are listed in Table 3. The prepositions used by primary rules in the table are omitted for brevity. See Appendix A.3 for a detailed list of these prepositions. Extraction is implemented by resorting to spaCy part-of-speech tagging and dependency parsing tools  [^3] For more details, see the spaCy documentation: https://spacy.io/usage/linguistic-features/#pos-tagging.
3.3 Representing
Figure 3: (a) Six-level hierarchical structure of PropNet. (b) An example of representing the sentence “A young man is playing an instrument in the garden”. Evolutionary nodes, developmental nodes, instance nodes and stamp nodes are labeled as red, orange, green and grey colors respectively.
As brain’s long-term memory stores information as a hierarchical network [18; 7; 37], PropNet represents a proposition using a six-level hierarchical network. Figure 3 (b) provides an example of how the sentence “A young man is playing an instrument in the garden” is represented. The network comprises three types of nodes: evolutionary nodes, developmental nodes and instance nodes which are represented as red, orange and green colors, respectively. In designing the network, we incorporate ideas similar to Nature Design and Nurture Beliefs , which are outlined in [52]. Evolutionary nodes correspond to Nature Design, which reflects human nature and is derived from heredity. They are denoted with a prefix “#”, like #action. Each dimension in the parsing phase has a corresponding evolutionary node. Developmental nodes correspond to Nurture Beliefs, which are learned from experience, and are named with a prefix “_”, such as _young. Each word in a sentence corresponds to a developmental node.
An evolutionary node cannot directly establish a connection with a developmental node. This connection is fulfilled by an instance node. Each evolutionary node points to at least one instance node. Each instance node can point to only one developmental node. They have the following three types:
  • Entity instance nodes are linked to developmental nodes whose lexical content has a hypernym “physical entity”, a relationship that can be ascertained using WordNet [43]. Entity instance nodes are denoted as ins_entity_id; for instance, ins_entity_17, which is linked to _man.
  • Concept instance nodes are linked to developmental nodes whose lexical content is not “physical entity”, such as adjectives, pronouns, and numbers. These nodes are denoted as ins_concept_id; for example, ins_concept_29, which is connected to _young.
  • Action instance nodes are linked to developmental nodes that contain verbal lexical content and are labeled as ins_action_id. As an example, ins_action_15 refers to _play.
Figure 3 (a) provides a more lucid depiction of the six-level hierarchical structure of PropNet. An action instance node functions as the root and triggers #action. It also points to a developmental node _play. This action instance node is unique because there is only one verb in a proposition. Edges from #action connect evolutionary nodes #subject, #object, #where, #aux_obj, #goal, #source, and #reason, each of which corresponds to a parsing dimension described in Section 3.2. These nodes will establish connections with instance nodes that are created for the words parsed from each dimension. For example, #subject points to ins_entity_17, which has a link with _man. At level 4, these instance nodes activate two evolutionary nodes, #attr and #part_of, which align with the parsing dimensions Attribute and Part_of respectively. Similarly, #attr and #part_of establish links with instance nodes created for the words parsed under Attribute and Part_of. For instance, ins_entity_17 points to #attr and #attr points to _young.
Strategy Rule Example
Merge Subjects Merge the instance nodes with the same subject into a single instance node. A cat sitting on sand looks up at the camera. $\rightarrow$ (a cat looks up at the camera, cat be sitting on sand) [Merge instance nodes of cat]
Link Clause Identifier to Action Stamp Node (ASN) Establish links between the instance node of identifier_ccomp, identifier_pcomp, xcomp_identifier, identifier_advcl and identifier_csubj in the MAIN and the ASN in the SUB. (MAIN: main proposition; SUB: subordinate proposition) She thinks this is a good idea. $\rightarrow$ (she thinks identifier_ccomp, this is a good idea) [Link the instance node of identifier_ccomp to the ASN invoked by is]
Link Action Stamp Nodes (ASNs) by Time Node MAIN Before SUB: Link the ASN of MAIN to #time_before, and then link #time_before to the ASN of SUB; MAIN After SUB: Link the ASN of SUB to #time_before, and then link #time_before to the ASN of MAIN; MAIN and SUB Occur Simultaneously: Link #time_same to the ASNs of MAIN and SUB. The boy washes his hands before he has lunch. $\rightarrow$ (the boy washes his hands, he has lunch) [ASN invoked by wash -> #time_before -> ASN invoked by has]
Table 4: Merging Strategies for A Prop2 Sentence
If during the parsing phase, no words are associated with a dimension, its corresponding evolutionary node will invoke an instance node that points to a default developmental node, _unknown. This ensures the network remains complete even when some information is missing. In Figure 3 (b), only #subject, #object and #where have valid instance nodes, which point to _man, _instrument and _garden, correspondingly. All other evolutionary nodes pointed to by #action are connected to instance nodes that point to _unknown, which is marked by a purple box in the figure.
There’s a fourth type of nodes, called stamp nodes, which can be distinguished by their grey color in Figure 3 (b). The primary role of stamp nodes is to distinguish different parts of the network representing a complex sentence that consists of several propositions. An action stamp node is activated by an action instance node and named as stamp_action_id. It points to all entity or concept instance nodes associated with that action, thereby grouping them together and differentiating them from the entity or concept instance nodes of other actions.
An entity stamp node is triggered by an entity instance node and named as stamp_entity_id. It points to all entity or concept instance nodes that describe the corresponding entity instance node. As shown by the blue box in the figure, stamp_entity_32 points to ins_concept_29, aggregating the modifier “young” to the entity “man”.
3.4 Merging
When the main proposition and the subordinate proposition of a Prop2 sentence are represented separately, we need to combine them to convey a complete and coherent meaning. We put forward three merging strategies, which are enumerated in Table 4. If the subordinate proposition is a relative clause, a different merging strategy is adopted, which is presented in Appendix A.2.
For a Prop3+ sentence, the merging process corresponds to the upward backtracking process of splitting, illustrated in Figure 2. The same strategies are applied as for Prop2. Appendix A.1 demonstrates the merging result of the sentence example in Figure 2.
3.5 Comparison of Two PropNets
Given the PropNets of a pair of sentences, a natural consideration is how to compare the differences between the two networks. However, since the PropNet of a complex sentence can be extensive, containing a large number of nodes and edges, directly comparing two PropNets poses a significant challenge due to computational complexity. A more viable computational approach involves comparing the disparities in the representations of the corresponding propositions between the two sentences, and then aggregating these disparities into an overall signature difference. Thus, we initially introduce the method for comparing the differences between two propositions, which is previously presented as pair type P1-. Next, based on this foundation, we discuss the difference-construction methods for more complex pair types P2 and P3+.
Figure 4: Framework of the comparison module. This module consists of two parts: difference vector computation and CART prediction. It is exemplified by the pair “The tall man is playing the delicate piano” and “The short man is playing the delicate guitar”. They differ at #action|#subject|#attr and #action|#object, which are marked with elements 2 and 1 at the corresponding positions in the difference vector. The codes 0, 1, 2 represent “identical”, “similar” and “different”, respectively. Note that the difference vector for P1- is padded with 0, doubling its size, before it enters the corresponding CART model.
3.5.1 Pair Type P1-
The comparison dimensions are formed by the Cartesian product of {#subject, #object, #where, #aux_obj, #goal, #source, #reason} and {$1$, #attr, #part_of}, further extended by {#action, #action|#subject|#where, #action|#object|#where}. For example, the dimension #action|#subject|#attr represents the attribute of an entity or concept that serves as the subject within the action. The reason of using the names of evolutionary nodes is this way of naming can reveals the route to locate the corresponding entity or concept instance nodes within the graph of PropNet. The explanations of all these 24 dimensions are provided in Appendix B.1.
By comparing the developmental nodes indicated by the dimensional instance nodes, we obtain a similarity code for each dimension. Concatenating these codes in sequence produces a difference vector $\mathbf{v_{1}}$. The similarity code primarily takes three values: 0 for identical, 1 for similar, 2 for different. A large language model (LLM) acts as the knowledge base to compute the similarity degree between two words. This strategy is word-level based, ignoring context-dependent meanings, e.g. polysemous words. Identification of the gloss that a word uses given a context is left in future work. Details on LLM choice and its prompts are explained in Appendix B.2. Additional similarity codes, and examples for each code are provided in Appendix B.3.
Figure 4 illustrates how to calculate a difference vector given the representations of two sentences “The tall man is playing the delicate piano” and “The short man is playing the delicate guitar”. Purple boxes identify the concept instance nodes under #action|#subject|#attr with two different developmental nodes _tall and _short. Blue boxes identify the entity instance nodes under #action|#object with two different developmental nodes _piano and _guitar. The instance nodes of other dimensions have exactly the same developmental nodes. Since _tall and _short are judged as different by LLM, #action|#subject|#attr contributes a code of 2 to the difference vector. As _piano and _guitar is judged as similar by LLM, #action|#object contributes a code of 1 to the difference vector.
3.5.2 Pair Type P2
Assume the first sentence s1 of the pair being compared can be split into two propositions, pp1 and pp2. If the second sentence s2 is Prop0 or Prop1, the following alignment process is applied. If two propositions share identical subjects and actions, they are considered to have a significant overlap. This criterion can be evaluated by examining the specific elements #action and #action|#subject of the difference vector. For a fair comparison, we align s2 to the proposition that exhibits a significant overlap with it. Subsequently, we utilize the difference vector associated with this alignment. This vector is then padded with 2 until it attains a length twice that of the original, resulting in the final output $\mathbf{v_{2}}$. Padding with 2 implies s2 lacks a corresponding component to compare with the other proposition of s1.
We illustrate the alignment process by comparing “I like apple” and “She likes orange and I like banana”. “She likes orange and I like banana” is split into “She likes orange” and “I like banana”. According to the order of appearance, we would compare “She likes orange” and “I like apple”, which is not a fair comparison. By implementing the alignment process, “I like banana” and “I like apple” are compared, while “She likes orange” has no counterpart for comparison.
If the second sentence s2 is Prop2, a similar alignment process is applied, but without padding. For example, in the pair “I like apple and she likes pineapple” and “She likes orange and I like banana”, the proposition “I like apple” aligns with “I like banana”, and “she likes pineapple” aligns with “she likes orange”.
3.5.3 Pair Type P3+
As a Prop3+ sentence can contain multiple propositions, we restrict our consideration to its first four propositions in the order of the leaf nodes in its splitting binary tree from left to right. Given a P3+ pair, suppose the first sentence $s_{1}$ and the second sentence $s_{2}$ have ordered proposition lists $L_{1}=[p_{1},p_{2},p_{3},p_{4}]$ and $L_{2}=[q_{1},q_{2},q_{3},q_{4}]$ respectively. Note that $p_{2},p_{3},p_{4}$ from $s_{1}$ and $q_{4}$ from $s_{2}$ might be empty strings. We calculate the Proposition Overlap Degree (POD) as the number of overlapping words between two propositions. $L_{2}$ is aligned with $L_{1}$ according to POD by the following steps:
  1. (1)
    The PODs between $p_{1}$ and each one of $L_{2}$ are calculated. Suppose $q_{1}^{(1)}$ from $L_{2}$ has the largest POD which is greater than a threshold value 1, then $q_{1}^{(1)}$ is moved to the front of $L_{2}$. Other elements in $L_{2}$ keep their orders invariant. After this rearrangement, $L_{2}$ is notated as $L_{2}=[q_{1}^{(1)},q_{2}^{(1)},q_{3}^{(1)},q_{4}^{(1)}]$. If the largest POD is no greater than 1, then $L_{2}$ keeps its original order, which means $L_{2}=[q_{1}^{(1)},q_{2}^{(1)},q_{3}^{(1)},q_{4}^{(1)}]=[q_{1},q_{2},q_{3},q_{4}]$.
  2. (2)
    Suppose $L_{2}^{(1)}=[q_{2}^{(1)},q_{3}^{(1)},q_{4}^{(1)}]$. Compute the PODs between $p_{2}$ and each one of $L_{2}^{(1)}$. Similarly, check if the largest POD is greater than 1, and rearrange $L_{2}^{(1)}$ to $[q_{2}^{(2)},q_{3}^{(2)},q_{4}^{(2)}]$.
  3. (3)
    Suppose $L_{2}^{(2)}=[q_{3}^{(2)},q_{4}^{(2)}]$. Compute the PODs between $p_{3}$ and each one of $L_{2}^{(2)}$. Check if the largest POD is greater than 1, and rearrange $L_{2}^{(2)}$ to $[q_{3}^{(3)},q_{4}^{(3)}]$. Finally, $s_{2}$ has a new ordered proposition list $[q_{1}^{(1)},q_{2}^{(2)},q_{3}^{(3)},q_{4}^{(3)}]$ which is aligned to $s_{1}$ according to POD.
After alignment, there are four P1- pairs: $p_{1}$ and $q_{1}^{(1)}$, $p_{2}$ and $q_{2}^{(2)}$, $p_{3}$ and $q_{3}^{(3)}$, $p_{4}$ and $q_{4}^{(3)}$. We compute a difference vector $\mathbf{v_{1}}$ for each of them using previous computing method for a P1- pair. The final difference vector for this P3+ pair is the concatenation of these four $\mathbf{v_{1}}$s. Therefore, the difference vector $\mathbf{v_{3}}$ for P3+ has a size that is twice the length of $\mathbf{v_{2}}$ for P2. If two propositions are both empty, $\mathbf{v_{1}}$ is replaced by a vector of the same size, where all elements are 0. If one proposition is empty and the other one is not, $\mathbf{v_{1}}$ is replaced by a vector of the same size, where all elements are 2.
Let’s exemplify the above processes, considering the pair: “Three men are jumping off a wall” and “Three young men run, jump, and kick off of a Coke machine”. The second sentence consists of three propositions: “three young men run”, “three young men jump”, and “three young men kick off of a Coke machine”. After the alignment process, its proposition list is reordered to “three young men jump”, “three young men run”, “three young men kick off of a Coke machine”. The first sentence is compared with “three young men jump”, resulting a difference vector $\mathbf{v_{1}}$. Since “three young men run” and “three young men kick off of a Coke machine” have no counterparts in the first sentence, we concatenate $\mathbf{v_{1}}$ with two vectors of the same length as $\mathbf{v_{1}}$, padded with 2 to signify the two sentences are different in this part. Finally, the resulting vector is padded with 0 until it reaches the required length, indicating the two sentences are the same in this default part.
4 Experiments
4.1 Comparison with Embedding Models
Model STS12 STS13 STS14 STS15 STS16 STS-B SICK-R Avg.
InferSent-GloVe (2017$\dagger$ $52.86$ $66.75$ $62.15$ $72.77$ $66.87$ $68.03$ $65.65$ $65.01$
USE (2018$\dagger$ $64.49$ $67.80$ $64.61$ $76.83$ $73.18$ $74.92$ $76.69$ $71.22$
SBERT (2019) $\dagger$ $70.97$ $76.53$ $73.19$ $79.09$ $74.30$ $77.03$ $72.91$ $74.89$
SimCSE-BERT (2021) $68.40$ $82.41$ $74.38$ $80.91$ $78.56$ $76.85$ $72.23$ $76.25$
AoE-BERT (2024) $75.26$ $85.61$ $80.64$ $86.36$ $82.51$ $85.64$ $80.99$ $82.43$
PropNet $\mathbf{47.01}$ $\mathbf{36.33}$ $\mathbf{40.74}$ $\mathbf{34.54}$ $\mathbf{12.06}$ $\mathbf{47.41}$ $\mathbf{55.35}$ $\mathbf{39.06}$
Table 5: Performance comparison with embedding models on STS tasks. We use Spearman’s rank correlation coefficient ($\rho$) multiplied by 100 as evaluation metric. Results of the models with $\dagger$ are from [47], others are from their original papers. There is a significant gap between PropNet and the latest embedding models.
To assess the performance gap between PropNet and mainstream approaches, PropNet is compared with widely used embedding methods on semantic textual similarity (STS) tasks.
Baselines. Given that our method is supervised, widely used supervised embedding methods are selected as baselines: InferSent [19], USE [17], SBERT [47], SimCSE [22], and AoE [39].
Datasets. We conduct experiments on three benchmark datasets: STS 2012-2016 [4; 5; 2; 1; 3], STS-B [16], and Sick-R [41].
Evaluation Metrics. We employ Spearman’s rank correlation coefficient ($\rho$) multiplied by 100 to measure the alignment between predicted and ground-truth scores.
Implementation Details. The experiment employs a comparison module which consists of two components, (1) the calculation of the difference vector and (2) prediction using Classification and Regression Trees (CART) [12]. Figure 4 presents the details of this module. After computing the difference vectors using the approach introduced in Section 3.5 for all pairs within the benchmark datasets, two CART models are trained on these vectors: one for short pairs P1-/P2 and another for long pairs P3+, because the sizes of the difference vectors are different. These models are denoted as CART(P1-/P2) and CART(P3+) respectively. We call the API provided by scikit-learn  [^4] API: https://scikit-learn.org/stable/modules/generated/sklearn.tree.DecisionTreeRegressor.html for training. Since STS 2012-2016 and STS-B share the same rules for ground scores, training is implemented only on the train set of STS-B to learn these rules. Similarly, for SICK-R, two CART models are trained on its training set. The random seed is fixed as 0, and the minimum number of samples at a leaf node is set to 10. Further training details are appended in Appendix C.1.
Note that the predictions for short pairs and long pairs in the test set are concatenated before computing the Spearman’s correlation.
Results. In Table 5, a significant gap is observed between PropNet and the latest embedding model, AoE, which scores above 80. This indicates the efficacy of big-data-trained embeddings in comprehending text semantics. This table also demonstrates that embedding models have undergone notable development over the last ten years, with scores ranging from 65.01 to 82.43. Since PropNet marks the inception of a completely new research direction, offering purely white-box and human-like sentence representation, we are confident that PropNet will improve in the future, considering that there is substantial room for enhancement in its current version, which is presented in the next section.
If we only focus on the scores of PropNet across all benchmarks, we observe a significant deviation, ranging from 12.06 on STS 2016 to 55.35 on SICK-R. The genre of STS 2016 is mainly forum question/answer and news, while SICK-R is image/video description. This genre-score discrepancy raises the question of whether PropNet has different representation abilities across different genres or if the comparison module fails for a specific genre. We investigate this factor in detail in the next section.
4.2 Investigation of Genres
In this section, we examine the performance of PropNet across different genres. The number of verbs is also considered as an important influencing factor. We anticipate that predicting accurately using our comparison module becomes more challenging for complex sentences.
Dataset. We utilize the STS-B test set  [^5] Approximately 8.7% of the STS-B records lack similarity scores due to spaCy dependency parsing errors. These records are excluded from the analysis. along with its predicted scores for analysis. The test set is split into three parts according to its field genre, which has three values: main-captions, main-news and main-forums. They represent image/video description, news and forum question/answer, respectively. Appendix C.2 provides more detailed explanations of these genres.
Another splitting dimension is verb number. The test set is divided into three parts according to pair types P1-, P2, and P3+. Generally, from P1- to P3+, each sentence in a pair contains an increasing number of verbs. Through intersection operation, pairs restricted to a specific genre and verb number can be selected.
Evaluation Metrics. Spearman’s rank correlation coefficient ($\rho$) multiplied by 100, is used to measure the alignment between predicted and ground truth scores for a specific subset of the STS-B test set. Moreover, the number of pairs in each subset, and their proportion relative to the entire test dataset are calculated to illustrate data distribution.
Total P1- P2 P3+
Count (%)   $\rho\times 100$ Count (%)  $\rho\times 100$ Count (%)  $\rho\times 100$ Count (%)  $\rho\times 100$
main-captions 616 (50%) 65.12 484 (39%) 70.41 120 (10%) 43.15 12 (1%) 28.09
main-news 411 (33%) 36.80 207 (17%) 40.37 124 (10%) 18.15 80 (6%) 20.22
main-forums 212 (17%) 19.98 91 (7%) 43.53 69 (6%) 26.70 52 (4%) -0.04
Total 1239 (100%) 47.41 782 (63%) 59.76 313 (25%) 26.11 144 (12%) 23.13
Table 6: Pair distributions and Spearman’s correlation scores corresponding to each genre and verb-number splitting. Count (%) represents the number of pairs and their percentage of the total number. The results indicate that: (1) PropNet performs significantly better on main-captions than the other two genres; (2) As the number of verbs increases, Spearman scores decline significantly across all genres; (3) main-news and main-forums contain relatively more P2 and P3+ pairs than main-caption, implying these two genres have more complex sentence structures.
Results. As shown in Table 6, PropNet scores 65.12 on main-captions, significantly higher than the 36.80 on main-news and 19.98 on main-forums. Even for the simplest pairs, P1-, PropNet attains only 40.37 on main-news and 43.53 on main-forums. This indicates that PropNet performs poorly on main-news and main-forums. The worst-performing pairs in P1-, whose prediction scores have the largest discrepancy with their ground scores, are analyzed to identify the underlying reasons.
For main-news, the lack of social common knowledge contributes to the failure of the comparison module for P1-. For instance, it cannot accurately measure the similarity between the pair "bitcoin haul" and "bitcoins" without knowledge of the terms "bitcoin" and "haul". Another reason is that the evolutionary graph of PropNet is too coarse to handle detailed comparisons. For example, in the sentence "At Northwest Medical Center of Washington County in Springdale , one child is in serious condition", the phrase "in serious condition" is incorrectly regarded as a description of #where. However, it should be linked to child’s #state, which is not included in current PropNet’s evolutionary graph.
For main-forums, PropNet currently lacks nodes to measure the likelihood of an action, such as modal verbs (e.g., "can", "should") and negation indicators (e.g., "not", "hardly"). For example, it fails in comparing “You should do it” and “You can do it”, or “It’s not a good idea” and “It’s a good question”. In addition, main-forums also suffers from the coarseness of the evolutionary graph. For instance, expressions like "It is up to you" cannot be represented properly.
Regarding verb number, Table 6 indicates that as the number of verbs increases, Spearman scores across all genres decrease substantially. The worst-performing pairs in P2 for main-news and main-forums are analyzed to identify the reason.
For main-news, the use of clear and concise language in its writing style sometimes causes difficulties for dependency parsing, leading to failed representations. For instance, consider the sentence "Coroner: Whitney Houston died in bathtub." PropNet cannot properly represent "Coroner: " as "Coroner claims".
For main-forums, sentences describing mental activities are often redundant, which complicates the alignment of propositions for computing the difference vector in the comparison module. For instance, in the pair "There are two things to consider" and "A couple things to consider", the phrase "there are" hinders proper alignment. A potential improvement could be simplifying "There are two things to consider" to "two things to consider" during comparison.
All the above case studies imply that there is a substantial improvement room in PropNet. Another point worthy of note is that, for main-news and main-forums, the internal proportions of P2 and P3+ are 49.6% (204/411) and 57.1% (121/212) respectively, which are significantly higher than 21.4% (132/616) of main-captions. This suggests that sentences in main-news and main-forums contain more complex structures.
4.3 Human Cognitive Differences in STS Tasks
Figure 5: (a) Mean Ground Scores. For main-captions, disparities in #action, #subject or #object tend to make the two sentences appear more dissimilar than #where and #other. Additionally, a difference in #action, #subject or #object is a sufficient condition for the non-equivalence of two propositions in human perception of sentence similarity. (b) Standard Deviation.
This section studies the cognitive process underlying how humans assign a score according to task instructions when given a sentence pair. Intuitively, different dimensions ought to contribute differently to the disparity between two sentences. In addition, for the same dimension, human perceptions of the disparity should vary to a certain extent. We aim to quantify these cognitive differences in this experiment.
Eight dimensions from Section 3.5.1 are considered: #action, #subject, #object, #where, #aux_obj, #goal, #source, and #reason. We aggregate #aux_obj, #goal, #source, and #reason into one dimension, labeled as #other.
Dataset. The P1- pairs of the STS-B dataset, whose sentences differ at only one dimension, are selected. Compared to the black-box nature of embedding models, PropNet significantly facilitates this selection process by simply checking if the corresponding element of difference vector is 2. For example, "A man is speaking" and "A man is cooking" is a valid pair for the #action dimension, since the difference vector has a value 2 at #action and 0 at all other positions. Only pairs from main-captions and main-news are used, since main-forums contains too few samples. In total, 133 valid pairs are collected for analysis. All pairs and their distribution are detailed in Appendix C.3.1.
Evaluation Metrics. The mean values of the ground-truth scores of the selected pairs within each dimension are calculated to reflect the human cognitive weight assigned to sentence similarity across the five dimensions. Additionally, the standard deviations of each dimension are provided to measure the cognitive deviations within a single dimension.
Results. In Figure 5 (a), for main-captions, the mean values of #action, #subject and #object are around 1.5, which are lower than those of #where and #other (around 2.5), with statistical significance at the 0.05 level. This indicates that disparities in #action, #subject or #object make the two sentences seem more different than the disparities in other two dimensions. For main-news, all dimensions have similar weights on difference perception, with no statistical significance.
Another discovery is that pairs with disparity within the dimension #action, #subject, or #object in both genres have ground scores below 2.5, which is the cut-off score for “not equivalent” and “roughly equivalent”. This implies that a difference in any of these three dimensions is a sufficient condition for the non-equivalence of two propositions in human perception of caption or news sentence similarity. The ground score rules are provided in Appendix C.3.3. Research regarding more complex sentence pairs, such as P2, is deferred to future work.
Figure 5 (b) illustrates the differences in standard deviation for manual scoring. No statistically significant human cognitive deviation is observed within any single dimension at the 0.05 level. Overall, main-captions exhibits a standard deviation of approximately 0.8. For main-news, the notable difference between #action and #other presented in the figure is only statistically significant at the 0.1 level. Future work aims to collect more samples to reevaluate cognitive deviations within individual dimensions. All hypothesis testing results are supplied in Appendix C.3.2.
5 Discussion
Multimodal Foundation. Instance nodes are necessary not only to process language, but also to empower PropNet with the potential to function as the backbone of understanding human’s vision and control. Imagine this: you see an apple on the table and stretch out your arm to pick it up. An instance node corresponding to this apple makes it possible to aggregate shape, color and name information together, distinguish this apple from other apples on the table, and also possible to track this apple and compute its relative positions to other objects when it is picked up.
Text Generation. If PropNet can represent an image or a video, the rules used in the splitting and parsing phases can be reversed to generate a meaningful sentence. For instance, given the subject "apple", the action "fall", and the action source "tree" perceived from an image or a video, PropNet can generate the sentence "apple fall from tree". This represents a novel approach to generate text, fundamentally distinct from the current mainstream multimodal large language models. It is completely transparent and is likely to be closer to the way humans generate language.
Long Text Understanding. Although in this work we mainly discuss the representation of a single sentence, it is possible to extend this method to a discourse or a conversation. Instance nodes enable PropNet to have potentials in addressing coreference resolution and representing contextual information effectively. Minor changes to current merging strategies can meet the need for linking the representations of each sentence into a unified network. This will provide novel perspectives for the research of long-text understanding in a white-box and human-like manner.
Cognitive Process Analysis. With the emergence of deep learning algorithms, the past decade has witnessed the generation of massive amounts of manually labeled data for training models. Therefore, cognitive processes underlying how humans label input data according to specific instructions, are recorded in these vast datasets. White-box and human-like approaches have the potential to illuminate the cognitive processes, which is exemplified by Section 4.3. It can significantly contribute to a deeper understanding of humanity, such as our behaviors, mentality, and intelligence.
6 Limitation
PropNet would fail to represent if there are errors from spaCy part-of-speech tagging and dependency parsing results. For instance, in the sentence “Blizzard wallops US Northeast, closes road”, the subject of “close” is identified as “wallop”.
7 Conclusion
In this paper, we have proposed a novel sentence representation network called PropNet. It can represent a complex sentence by a hierarchical network with interpretable and transparent structure. Through case analysis, several feasible optimization directions are identified to narrow the gap with SOTA models in STS tasks. PropNet paves the way for analyzing and understanding the human cognitive processes underlying STS benchmarks.
Appendix A PropNet Construction
A.1 Cases
Figure 6 shows the PropNet representation of a sentence with three propositions. The instance nodes pointing to _unknown are all omitted, as well as their evolutionary nodes, in order to provide a clean view. From this representation, we can form a picture of what happens in our mind.
A more complex sentence with five propositions is presented in Figure 7. Its splitting and merging processes are explained in Figure 2. Notably, the subject of the last proposition is incorrect. The correct result should be “man takes it out on the water”. A more robust strategy of subject identification in the splitting phase will be added in future work.
Figure 6: PropNet for “A man is sitting in a chair wearing a cloak and holding a stick.” It merges the representations of three propositions: “a man is sitting in a chair”, “man be wearing a cloak”, and “man be holding a stick”, each enclosed by separate grey frames.
Figure 7: PropNet for “A black and white dog jumping in the air to catch a Frisbee when a man is getting his boat clean to take it out on the water.” It merges the representations of five propositions: “a black and white dog jumping in the air to identifier_advcl”, “dog catch a Frisbee”, “a man is getting identifier_ccomp to identifier_advcl”, “his boat clean”, and “boat take it out on the water”, each enclosed within five grey frames.
A.2 Relative Clause
As a relative clause is more complex than other types of clauses, additional rules are proposed to address its splitting and merging. Currently, only relative pronouns that, which, who and relative adverbs when, where are considered for a Prop2 sentence with a relative clause. Table 7 shows the splitting and merging rules with examples. The merging results of these examples are shown in Figure 8, Figure 9 and Figure 10, respectively. It should be noted that the action stamp nodes of main proposition and subordinate proposition are linked by the evolutionary node #time_same. This merging rule is not included in the table.
Table 7: Splitting and Merging Rules for a Prop2 Sentence with a Relative Clause.
Type Splitting Rule Merging Rule Example
Relative Pronouns If t.lemma_ in (that, which, who) and t.dep_ in (nsubj, dobj), replace t by the antecedent in subordinate proposition. Merge the two instance nodes of antecedent. The book that I borrowed from the library is interesting. -> (the book is interesting, I borrowed book from the library) [Merge instance nodes of book.]
Relative Adverbs If t.lemma_ in (when, where) and t.dep_ in (advmod), keep t in subordinate proposition. Link instance node of antecedent to action stamp node of subordinate proposition. I like the place where I live in the street. -> (I like the place, where I live in the street) [Link instance node of place to action stamp node invoked by live]
No Relatives No processing in subordinate proposition. Link instance node of antecedent to action stamp node of subordinate proposition. The book I borrowed from the library is interesting. -> (the book is interesting, I borrowed from the library) [Link instance node of book to action stamp node invoked by borrowed]
Figure 8: A relative pronoun example “The book that I borrowed from the library is interesting.” It is split into two propositions: “the book is interesting”, “I borrowed book from the library”. The instance nodes of book are merged into ins_entity_17.
Figure 9: A relative adverb example “I like the place where I live in the street.” It is split into two propositions: “I like the place”, “where I live in the street”. The instance node of place, ins_entity_19, is linked to the action stamp node stamp_action_41, invoked by live.
Figure 10: A no-relative example “The book I borrowed from the library is interesting.” It is split into two propositions: “the book is interesting”, “I borrowed from the library”. The instance node of book, ins_entity_17, is linked to the action stamp node stamp_action_41, invoked by borrow.
A.3 Parsing
This section provides a detailed explanation of the primary parsing rules. For Action, Subject, Object, Attribute and Part_of, the parsing primarily relies on token.dep_ and token.pos_. For Where, Auxiliary_Object, Goal, Reason and Source, the parsing system depends on detecting and classifying token’s prepositions. The classification is based on the definitions of prepositions from the Merriam-Webster Dictionary. We acknowledge that current classification remains course and occasionally inaccurate. For example, in “shoot at an elephant”, “elephant” is classified as Where. However, this Where should be interpreted as the action goal in conjunction with the meaning of "shoot". This issue can be addressed by adding more evolutionary nodes. Additionally, the temporal attributes of an action, such as its timing, are temporarily assigned to Where. More sophisticated space-time modules are planned for development in future research.
The complete list of preposition categories is as follows:
  • Where: “on”, “in”, “inside”, “through”, “over”, “around”, “down”, “at”, “near”, “along”, “outside”, “past”, “across”, “during”, “after”, “before”, “while”, “whilst”, “off”, “amid”, “behind”.
  • Auxiliary_Object: “with”, “by”, “about”, “like”, “as”.
  • Goal: “into”, “to”, “onto”, “towards”, “against”
  • Reason: “for”, “due”.
  • Source: “from”.
Appendix B Comparing Two Sentences
B.1 Comparison Dimensions
Table 8 explains the comparison dimensions to compute a difference vector given two sentence representations. Be aware that since instance nodes themselves have no meanings, what are really compared are the corresponding developmental nodes.
Dimension Explanation
#action Find and compare all the instance nodes that represent actions.
#action|#subject Find and compare all the instance nodes that represent the subject of an action.
#action|#subject|#attr Find and compare all the instance nodes that belong to attributes of an entity or concept that serves as the subject of the action.
#action|#subject|#part_of Find and compare all the instance nodes that represent parts of an entity or concept that serves as the subject of the action.
#action|#subject|#where Find and compare all the instance nodes that represent the location of the subject of the action.
#action|#object Find and compare all the instance nodes that represent the object of an action.
#action|#object|#attr Find and compare all the instance nodes that belong to attributes of an entity or concept that serves as the object of the action.
#action|#object|#part_of Find and compare all the instance nodes that represent parts of an entity or concept that serves as the object of the action.
#action|#object|#where Find and compare all the instance nodes that represent the location of the object of the action.
#action|#aux_obj Find and compare all the instance nodes that represent auxiliary objects of an action.
#action|#aux_obj|#attr Find and compare all the instance nodes that belong to attributes of an entity or concept that serves as the auxiliary object of the action.
#action|#aux_obj|#part_of Find and compare all the instance nodes that represent parts of an entity or concept that serves as the auxiliary object of the action.
#action|#where Find and compare all the instance nodes that represent the location of the action.
#action|#where|#attr Find and compare all the instance nodes that belong to attributes of the location of the action.
#action|#where|#part_of Find and compare all the instance nodes that represent parts of the location of the action.
#action|#goal Find and compare all the instance nodes that represent the goal of the action.
#action|#goal|#attr Find and compare all the instance nodes that belong to attributes of the goal of the action.
#action|#goal|#part_of Find and compare all the instance nodes that represent parts of the goal of the action.
#action|#source Find and compare all the instance nodes that represent the source of the action.
#action|#source|#attr Find and compare all the instance nodes that belong to attributes of the source of the action.
#action|#source|#part_of Find and compare all the instance nodes that represent parts of the source of the action.
#action|#reason Find and compare all the instance nodes that represent the reason for the action.
#action|#reason|#attr Find and compare all the instance nodes that belong to attributes of the reason for the action.
#action|#reason|#part_of Find and compare all the instance nodes that represent parts of the reason for the action.
Table 8: Comparison Dimensions for Computing A Difference Vector.
B.2 Similarity between Two Words
Given two words, we use Doubao  [^6] https://team.doubao.com/en/direction/llm to compute a similarity score ranging from 0 to 1. Doubao has an extremely low inference cost, while offering top-tier inference performance and an easy-to-use API for calling. The quality of the output scores are validated by manual inspection during development. Since this task is straightforward, other LLM models like ChatGPT or Llama should work as well.
For a verb pair, its prompt is "From physical action perspective, return similarity score (0~1) between two verbs: {word_text1} and {word_text2}. Only return the score".
For other pairs, its prompt is "From semantic perspective, return similarity score (0~1) between two words: {word_text1} and {word_text2}. Only return the score".
Code Meaning Calculation Rules Examples
0 Completely or Almost the Same (1) Two words have the same text; (2) LLM similarity score >= 0.7; (3) binary group similarity score >= 0.7 (1) "guitar" vs "guitar"; (2) "cat" vs "kitten" (3) ["red", "cat"] vs ["pink", "cat"]
1 Similar (1) 0.2 <= LLM similarity score < 0.7; (2) One side is "unknown"; (3) 0.2 <= binary group similarity score < 0.7 (1) "guitar" vs "piano"; (2) "guitar" vs "unknown"; "unknown" vs ["guitar", "piano"] (3) ["short", "man"] vs ["tall", "man"]
2 Different (1) LLM similarity score < 0.2 (2) binary group similarity score < 0.2 (1) "tall" vs "short" (2) ["short", "man"] vs ["big", "cat"]
3 Long and Equal Length Two group words with equal length which is greater than 2. ["violin", "guitar", "piano"] vs ["car", "bus", "truck"]
4 Different Length (1) Single word and a group of words; (2) Two group words with different length. (1) "guitar" vs ["violin", "piano"] (2) ["guitar", "piano"] vs ["car", "bus", "truck"]
Table 9: Dimensional Similarity Code.
B.3 Dimensional Similarity Code
Given a comparison dimension, it’s common that both sentences have only one developmental node. For example, considering this pair "I like orange" and "she hates apple", #action is "like" and "hates", #action|#subject is "I" and "she", and #action|#object is "orange" and "apple". But exceptions do exist that the developmental nodes might be more than one. For instance, in the representation of "I like orange and apple", #action|#object has two developmental nodes _orange and _apple.
We define Dimensional Similarity Code (DSC) as an indicator for assessing similarity between two sentence representations under a specific comparison dimension. Three scenarios are considered: (1) comparing two words, (2) comparing a single word with a group of words, and (3) comparing two groups of words. The DSC codes and their corresponding computing rules are explained in Table 9.
Given two words, a code of 0 means two words are completely or almost the same. A code of 1 means they are similar, sharing some important details. Note that when we compare "unknown" with one single word or a group of words, they are regarded as similar. For example, in the comparison of "I cut an onion with a knife" and "I cut an onion", the dimension #action|#aux_obj has "knife" for the first sentence and "unknown" for the second sentence, and this comparison receives a code of 1. A code of 2 means the two words are completely different. We manually check word pair cases along with their LLM scores to determine proper thresholds, best matching the DSC code standard. If LLM similarity score is no less than 0.7, a code of 0 is assigned. If LLM similarity score is less than 0.2, a code of 2 is assigned. Otherwise, a code of 1 is assigned.
When comparing one single word with a group of words, or two groups of words of different lengths, we label this scenario as "different length" with a code of 4. An exception occurs when the single word is "unknown", in which case it receives a code of 1. Further refinement of similarity discrimination, incorporating semantic analysis, is left for future work.
Given two groups of words with equal length greater than 2, we label this situation as "long and equal length" and assign it a code of 3. More refined similarity discrimination, combined with semantics, is deferred to future work.
Finally, given two groups of words with an equal length 2, we consider a binary group similarity score which is calculated by the following steps:
  1. (1)
    Initialize score as 0.
  2. (2)
    Directly calculate the intersection between two groups of words and obtain the length of the intersection.
  3. (3)
    Add the length of the intersection to the score variable.
  4. (4)
    Remove the intersecting parts from each word group.
  5. (5)
    If each group contains only one remaining word, calculate the similarity score between these two words and add it to score. Otherwise do nothing.
  6. (6)
    Divide score by 2 to obtain the binary group similarity score.
This process is exemplified step by step by comparing ["apple", "banana"] and ["apple", "orange"] as follows:
  1. (1)
    Initialize score as 0.
  2. (2)
    Intersection is ["apple"] and its length is 1.
  3. (3)
    score = 1 + score. Now score is 1.
  4. (4)
    Remove "apple" from both groups. Now they become ["banana"] and ["orange"].
  5. (5)
    Since each group has only one word left, LLM is used to calculate the similarity score which outputs a 0.2. Now score is 1.2.
  6. (6)
    Binary group similarity score = score / 2 = 0.6
As the final score lies between 0.2 and 0.7, a DSC code 1 (similar) is assigned for this binary-word groups comparison.
Appendix C Experiments
C.1 CART Training Details for STS Tasks
For the STS-B dataset, we tune only one parameter for CART training, the minimum number of samples at a leaf node (short as min_samples_leaf), on its train.jsonl (5749 samples) and validation.jsonl (1500 samples). The default values are adopted for other CART parameters, since finetuning models for higher Spearman’s score is not a key point in our work. What we place high weight on, is finding the weaknesses of PropNet and its opitimization directions in future research.
The SICK-R dataset contains a field named SemEval_set, indicating the train/dev/test split. We use the training split (4439 samples) and dev split (495 samples) for tuning min_samples_leaf.
CART Training is supported by the Scikit-Learn APIs, whose version is 1.6.0.
C.2 Genres
STS-B dataset has a field genre with three values, which are explained as follows:
  • main-captions: image/video description, focusing on the main action or event in a video or an image. For example, "A woman is spreading mustard on a bread roll." The words have strong relations with physical concepts, as defined in [53].
  • main-news: news, for instance, "Stones add second Hyde Park concert". The words have strong relations with social concepts.
  • main-forums: forum question or answer, for instance, "This is a part answer to your question". The words appearing in such a sentence are mainly mental, as introduced in [53].
Based on the action types introduced in [52], a proposition of main-captions, main-news and main-forums correspond to physical, mental and social types, respectively.
C.3 Cognitive Differences in STS Tasks
C.3.1 Experimental Samples
Due to insufficient pairs for #aux_obj, #goal, #source, and #reason, these four dimensions are aggregated into one, labeled as #other. Valid pairs from STS 2012-2016 are added for the #other dimension. The experiment’s results are not affected by this supplement, since the scoring criteria remain consistent across STS 2012-2016 and STS-B datasets.
All pair examples selected for main-captions and main-news in this experiment are listed in Table LABEL:tab:exp3_caption_examples and Table LABEL:tab:exp3_news_examples respectively. Pair distributions, ground score means and standard deviations for main-captions and main-news are listed in Table 10 and Table 11 respectively.
Dimension Count Proportion (%) Mean STD
#action 24 27.91 1.40 0.83
#subject 12 13.95 1.55 0.83
#object 25 29.07 1.70 0.56
#where 21 24.42 2.71 0.85
#other 4 4.65 2.53 0.80
Table 10: Pair statistics for main-captions. The total number of pairs is 86.
Dimension Count Proportion (%) Mean STD
#action 6 12.77 1.83 0.48
#subject 10 21.28 1.99 0.77
#object 12 25.53 2.45 0.80
#where 9 19.15 2.60 0.95
#other 10 21.28 2.32 1.36
Table 11: Pair statistics for main-news. The total number of pairs is 47.
C.3.2 Statistical Testing
Mann-Whitney U Test is performed to detect the statistical difference in ground score means between any two dimensions, as the ground scores are not normally distributed for all dimensions using Shapiro-Wilk Test. The significance level is set to 0.05. Results for main-captions are reported in Table 12, indicating #action, #subject and #object are significantly different from #where and #other. Although #subject and #other do not reach the significance level of 0.05, they are close to the significance level of 0.1. The situation for #object and #other is also similar. Results for main-news are reported in Table 13, revealing that there is no statistical significance.
The Levene’s Test is conducted to assess the statistical difference in the ground score standard deviations between all pairs of dimensions. The significance level is set to 0.05. The results for main-captions and main-news are reported in Table 14 and Table 15, respectively. We do not find any statistically significant differences.
Dimension1 Dimension2 U-statistic P-value Statistical Significance (0.05)
#action #subject 121.5 0.4595 No
#action #object 217.5 0.1000 No
#action #where 73.0 0.0000 Yes
#action #other 16.5 0.0413 Yes
#subject #object 126.5 0.4545 No
#subject #where 47.0 0.0032 Yes
#subject #other 10.5 0.1141 No
#object #where 84.5 0.0001 Yes
#object #other 21.5 0.0752 No
#where #other 48.0 0.6815 No
Table 12: Results of the Mann-Whitney U Test for comparing main-captions dimensions.
Dimension1 Dimension2 U-statistic P-value Statistical Significance (0.05)
#action #subject 23.5 0.5114 No
#action #object 20.5 0.1570 No
#action #where 14.0 0.1368 No
#action #other 19.0 0.2526 No
#subject #object 43.0 0.2747 No
#subject #where 31.0 0.2682 No
#subject #other 36.0 0.3049 No
#object #where 49.5 0.7755 No
#object #other 58.5 0.9472 No
#where #other 47.5 0.8698 No
Table 13: Results of the Mann-Whitney U Test for comparing main-news dimensions.
Dimension1 Dimension2 Statistic P-value Statistical Significance (0.05)
#action #subject 0.0 0.9566 No
#action #object 1.5 0.2269 No
#action #where 0.0 0.9599 No
#action #other 0.0 0.9543 No
#subject #object 1.5 0.2234 No
#subject #where 0.0 0.9935 No
#subject #other 0.0 0.9291 No
#object #where 1.4 0.2443 No
#object #other 0.6 0.4614 No
#where #other 0.0 0.9378 No
Table 14: Results of the Levene’s Test for comparing main-captions dimensions.
Dimension1 Dimension2 Statistic P-value Statistical Significance (0.05)
#action #subject 2.0 0.1793 No
#action #object 3.1 0.0986 No
#action #where 3.4 0.0868 No
#action #other 3.2 0.0967 No
#subject #object 0.1 0.7885 No
#subject #where 0.7 0.4097 No
#subject #other 1.6 0.2186 No
#object #where 0.5 0.5006 No
#object #other 1.5 0.2343 No
#where #other 0.4 0.5562 No
Table 15: Results of the Levene’s Test for comparing main-news dimensions.
C.3.3 Ground Score Rules
The ground score rules for STS-B and STS 2012-2016 are provided below, following the original papers [16, 5].
  • Score 5: The two sentences are completely equivalent, as they mean the same thing.
  • Score 4: The two sentences are mostly equivalent, but some unimportant details differ.
  • Score 3: The two sentences are roughly equivalent, but some important information differs/missing.
  • Score 2: The two sentences are not equivalent, but share some details.
  • Score 1: The two sentences are not equivalent, but are on the same topic.
  • Score 0: The two sentences are completely dissimilar.
Table 16: Sentence pair examples in main-captions with only one different dimension.
Dimension Sentence1 Sentence2 Ground Score Data Source
#action A man is cycling A man is talking 0.6 STS-B
#action A man is speaking A man is cooking 0.8 STS-B
#action A man is praying A man is dancing 0.75 STS-B
#action A man is dancing A man is thinking 1.2 STS-B
#action people walk home People waiting 1.6 STS-B
#action A man is speaking A man is spitting 0.636 STS-B
#action A man is dancing A man is singing 1.25 STS-B
#action A man is running A man is singing 1.25 STS-B
#action A man is speaking A man is dancing 1.2 STS-B
#action A man is dancing A man is speaking 1.2 STS-B
#action Two dogs fighting in the snow Two dogs standing in the snow 2.4 STS-B
#action two dogs running in the snow A dog laying in the snow 2.2 STS-B
#action The people are leaving the airplane The people are entering the plane 2.2 STS-B
#action Two men are sitting in the room Two men are standing in a room 3.0 STS-B
#action A woman is writing A woman is swimming 0.5 STS-B
#action A man is levitating A man is talking 0.8 STS-B
#action A woman is slicing up some meat A woman is breading some meat 2.25 STS-B
#action A man is praying A man is running 0.727 STS-B
#action A man is singing A man is dancing 0.5 STS-B
#action A man is running A man is mooing 1.0 STS-B
#action A man is dangling a mouse near a snake A man is feeding a mice to a snake 3.2 STS-B
#action A animal is eating The animal is hopping 0.4 STS-B
#action A group of people eat at a table outside Group of elderly people sitting around a table 2.8 STS-B
#action A man and a woman laughing A man and a woman kiss 1.2 STS-B
#subject The ballerina is dancing A man is dancing 1.75 STS-B
#subject Men are playing soccer Two teams play soccer 3.0 STS-B
#subject A slow lori walks around A animal is walking around 2.8 STS-B
#subject The lamb is looking at the camera A cat looking at the camera 0.8 STS-B
#subject A skateboarder jumps off the stairs A dog jumps off the stairs 0.8 STS-B
#subject Raccoons are eating A man is eating 1.5 STS-B
#subject A Woman is eating A animal is eating 1.6 STS-B
#subject A boy and a girl is dancing in the rain A man and woman is dancing in the rain 2.6 STS-B
#subject A black and white cow standing in a grassy field A blue jay standing in a grassy field 0.6 STS-B
#subject Old green bottle sitting on a table Three men in suits sitting at a table 0.4 STS-B
#subject Long - haired black dog standing in grassy field A blue jay standing in a grassy field 1.4 STS-B
#subject Three dogs racing on a dirt track Cars racing on a dirt track 1.4 STS-B
#object A woman is cutting onions A woman is cutting tofu 1.8 STS-B
#object A man is slicing a tomato A man is slicing a bun 2.0 STS-B
#object A woman is cutting tofu A woman is cutting an onion 2.4 STS-B
#object A woman is cutting some fish A woman is cutting tofu 2.2 STS-B
#object A woman is slicing ginger A woman is cutting potatoes 1.6 STS-B
#object A group of kids are having a jumping contest A group of kids are having a sleepover 1.2 STS-B
#object The dog is chasing the geese One dog is chasing the other 1.6 STS-B
#object A man is playing soccer A man is playing flute 1.0 STS-B
#object A girl is riding a horse A girl is riding a bicycle 1.917 STS-B
#object A man is playing a basketball A man is playing a piano 1.4 STS-B
#object A man is playing an electronic keyboard A man is playing a flute 1.2 STS-B
#object A guy is playing hackysack A man is playing a key - board 1.0 STS-B
#object A woman is slicing a potato A woman is slicing carrot 2.5 STS-B
#object A man is cutting up a potato A man is cutting up carrots 2.375 STS-B
#object The gate is blue The gate is yellow 1.6
#object A man is slicing a bun A man is slicing an onion 2.4 STS-B
#object A man is cutting up a potato A man is cutting up carrots 2.375 STS-B
#object A woman is peeling a potato A woman is peeling an apple 2.0 STS-B
#object The men are playing cricket The men are playing basketball 2.2 STS-B
#object A woman is doing weight exercises A woman is doing her hair 0.5 STS-B
#object The man is slicing a fish open A man is slicing a potato 2.2 STS-B
#object A woman is cutting an apple A woman is cutting potato 1.8 STS-B
#object A woman is cutting some flowers A woman is cutting broccoli 1.0 STS-B
#object A woman is cutting potatoes A woman is slicing carrots 1.2 STS-B
#object A woman is chopping garlic A woman slices a fish 1.0 STS-B
#where A yellow bird is eating fruit on a wire grate A yellow bird eating fruit on a bird feeder 2.8 STS-B
#where Two dogs play in the grass Two dogs playing in the snow 2.8 STS-B
#where A brown and white dog is running through the snow a brown and white dog is running on the grass 2.4 STS-B
#where Three children playing in snow Three children playing in hay 2.4 STS-B
#where Two dogs playing in snow Two dogs playing in grass 2.4 STS-B
#where A jockey riding a horse in a pen A jockey rides a horse at a gallop 3.6 STS-B
#where A red and white plane flying on a sunny day Red and white plane flying through the air 4.2 STS-B
#where Two women sitting in lawn chairs Two women are sitting in a cafe 2.6 STS-B
#where A white cat laying on an office chair A white cat laying on a sheet 2.4 STS-B
#where Two girls walking in the ocean two girls walking in the street 1.8 STS-B
#where Three dogs are playing in the white snow Two dogs are playing in the grass 1.6 STS-B
#where A man and woman are driving down the street in a jeep A man and woman are driving down the road in an open air vehicle 4.0 STS-B
#where Two men are fighting in a cow pasture Two men are fighting in a cattle pen 4.0 STS-B
#where People sitting on the porch People sitting on a couch 1.4 STS-B
#where A white jeep parked in front of a store A white Jeep parked on a street 3.5 STS-B
#where The two dogs are running through the grass Three dogs run in the snow 2.2 STS-B
#where A dog runs through the grass A dog runs through the snow 2.2 STS-B
#where A black dog running in the snow A black dog running on a beach 1.8 STS-B
#where a man walks two dogs on leashes down the street A man walks two dogs in the city 4.2 STS-B
#where A dog is running in the snow Two dogs running in the dirt 2.2 STS-B
#where A brown dog is running through green grass A brown dog is running though a river 2.4 STS-B
#other The dog is running with food in his mouth The dog is running with a yellow ball in his mouth 2.2 STS-B
#other A baby tiger is playing with a ball A baby is playing with a doll 1.6 STS-16
#other A man is running with a bus A man runs with a truck 2.5 STS-16
#other A cat is rubbing against baby’s face a cat is rubbing against a baby 3.8 STS-12
Table 17: Sentence pair examples in main-news with only one different dimension.
Dimension Sentence1 Sentence2 Ground Score Data Source
#action Saudi Arabia gets a seat at the UN Security Council Saudi Arabia rejects seat on UN Security Council 2.8 STS-B
#action Romney wins Florida Republican primary Romney eyes US Republican primary endgame 1.8 STS-B
#action Hong Kong stocks close down 0.28 % Hong Kong stocks open 0.62 pct higher 1.6 STS-B
#action French train derails south of Paris French train passengers tell of crash ordeal 1.2 STS-B
#action Indian stocks open lower Indian stocks close lower 1.8 STS-B
#action FAA continues ban on US flights to Tel Aviv FAA lifts ban on U.S. flights to Tel Aviv 1.8 STS-B
#subject It has a margin of error of plus or minus 4 percentage points That poll had 712 likely voters and sampling error of plus or minus 3.7 percentage points 2.5 STS-B
#subject 20 killed in bomb attack at Pakistani base 8 soldiers killed in bomb attack in NW Pakistan 2.0 STS-B
#subject Pak religious body endorses underage marriage CII endorses underage marriage 3.0 STS-B
#subject At least 15 killed in Nigeria church attack At least 12 killed in Nigeria church bombing 2.6 STS-B
#subject Brooks pleads not guilty to hacking charges Dave Lee Travis pleads not guilty to all charges 1.4 STS-B
#subject tirana is the capital of abuja is the capital of 1.0 STS-B
#subject Santorum’s 3 - year - old daughter hospitalized ’Kony 2012’ director hospitalized 0.6 STS-B
#subject 7 killed in attacks in Iraq 27 killed in attacks across Iraq 2.0 STS-B
#subject Six Australians killed in Laos plane crash Dozens killed in Laos plane crash 3.0 STS-B
#subject Philip leaves hospital after 11 days Mandela leaves hospital after 10 days 1.8 STS-B
#object US drone strike kills 11 in Pakistan US drone kills 16 in Pakistan 3.2 STS-B
#object Suicide attack kills eight in Baghdad Suicide attacks kill 24 people in Baghdad 2.4 STS-B
#object US drone strike ’kills 4 militants in Pakistan’ US drone strike kills 10 in Pakistan 3.8 STS-B
#object US drone strike kills four in North Waziristan U.S. drone strike kills 10 in northwest Pakistan: officials 1.8 STS-B
#object Indonesian president to visit UK Indonesian president to visit Australia 1.4 STS-B
#object 6.0-magnitude quake hits northern Italy: USGS 5.9-magnitude quake hits Sunda Strait, Indonesia: USGS 1.6 STS-B
#object 6.3-magnitude earthquake hits Taiwan 7.7-magnitude earthquake hits SW Pakistan 1.2 STS-B
#object Roadside bombs kill 5 in Afghanistan Roadside bomb kills 3 policemen in Afghanistan 3.4 STS-B
#object US drone strike kills three in northwest Pakistan US drone strike kills seven in North Waziristan 2.4 STS-B
#object Iraq violence kills 11 Iraq violence kills seven 2.2
#object Strong earthquake in western China kills 47 people Strong earthquake in western China kills at least 75 3.0 STS-B
#object Bomb attacks kill 20 in Baghdad’s Christian areas Bombs Kill 35 in Baghdad Christian Area 3.0 STS-B
#where Pakistan blocks Twitter over anti - Islamic material Pakistan blocks Twitter over ’ blasphemy ’ 3.8 STS-B
#where Shenzhen stock indices close higher Monday Shenzhen stock indices close lower – Oct. 31 1.8 STS-B
#where Two NATO soldiers killed in Afghanistan NATO soldier killed in Afghan attack 4.0 STS-B
#where NATO Soldier Killed In Afghan Attack NATO soldier killed in S. Afghanistan 3.8 STS-B
#where Two NATO soldiers killed in Afghanistan NATO Soldier Killed in Afghan Blast 3.6 STS-B
#where ’ Scores of bodies ’ found in Syria Eight more bodies found on ship 1.4 STS-B
#where Musharraf arrested in Lal Masjid case Musharraf arrested in Pakistan 3.0 STS-B
#where Hushen 300 Index closes higher – Oct. 14 Hushen 300 Index closes lower – March 12 2.0 STS-B
#where Seven peacekeepers killed in Sudan’s Darfur Peacekeeper killed in Abyei clash 1.6 STS-B
#where Powerful 7.6 quake strikes off Solomons Powerful 6.9 quake strikes off California coast 2.2 STS-B
#other 5 Things to Know About the Sochi Olympics 7 Things to Know About Ethanol 0.0 STS-B
#other van der merwe sentenced geiges to a total of 13 years imprisonment van der merwe suspended geiges’ sentence to 5 years imprisonment 2.6 STS-B
#other Death toll in Nigeria police attack rises to 30 Death toll in Kenya bus attack rises to six 1.4
#other Residents return to Texas blast site Residents return to Fallujah 0.0 STS-B
#other China yuan strengthens to 6.2689 against USD China yuan strengthens to new high against USD 3.0 STS-B
#other Death toll from Philippine earthquake rises to 185 Death toll from Philippines quake rises to 144 3.6 STS-B
#other Stars pay tribute to Cory Monteith Stars pay tribute to James Garner 2.2 STS-B
#other Death toll from Egypt protests rises to 49 Death toll from Egypt violence rises to 638 3.2 STS-B
#other Newark mayor saves neighbor from fire Newark mayor rescues neighbor from burning house 4.2 STS-B
#other 10 Things to Know for Wednesday 10 Things to Know for Today 3.0 STS-B